# Jaime Hernández-Lasheras

### v1

@@ 4 Discussion
The experiments presented here apply an approach which has not been used before for the design and extension of a HFR system. The NR is validated to give realistic simulations and the innovation distribution of the pseudo-observations present a similar distribution to the real data. The results obtained in the OSSE framework in terms of error reduction and correlation increment are of the same order as of the ones obtained in the previous OSE work. For surface currents, there is a significant difference between the innovations obtained in OSE and OSSE, however we have seen that this difference is not systematic and depends on the analyzed period (Figure 6). Furthermore, when assimilating HFR, the error reduction and the increment in the correlation are also of the order of the one achieved in the OSE.
@@
We have evaluated the impact on the surface currents in a wider area than that covered by the antennas, taking advantage of the full ocean state knowledge provided by the OSSE framework. It draws attention that the assimilation of only generic observing sources cannot correct the circulation in the area in the OSSE. In the previous OSE the assimilation of generic sources led to a good improvement compared to the CR. Even though the model’s response to SLA, SST and Argo observations is very similar to that obtained with real observations, the circulation seems to have a highly ageostrophic component. Therefore, the correction of SSH and density fields is not able to correct the surface currents in this region. This enhances the need of high-resolution surface current observations in coastal areas where satellite-derived geostrophic currents tend to fail (Vignudelli et al., 2019).
@@
The extension of the HFR system seems to be useful under certain conditions. The new antennas provide a moderate effect when NR and CR simulations reproduce similar circulation patterns. However, when both simulations present more differences, especially in the western side of the channel, having surface current observations of the full channel is key to improve the model dynamics. This can be interpreted as real situations in which the model is unable to reproduce the dynamics observed by the HFR system.
@@
The observation error for the HFR has been considered the same in all the domain.It could be supposed that in areas covered by three antennas more radial observations will be used to generate a total observation, reducing the expected error. This approach, using a spatial variable error depending on the number on antennas that cover an area has been explored by some authors (Vandenbulcke et al., 2017). The improvement of the observations error, including correlated errors remains as a potential aspect to explore in future studies. Besides, the generation of all pseudo-observations could be made more realistic by generating the random noise based on a spatial structure, using a given correlation length.
@@
We used a Lagrangian approach to evaluate the impact of data assimilation on the surface transport. The Lagrangian techniques, such as FSLE, have been increasingly being used in the last years. However, the effects of using a model field sequentially corrected through DA remains still poorly studied. As particles are advected, the DA simulation’s discontinuities might impact the trajectories of the particles and the following computation of FSLE. Here we showed that simulations using DA behave in a similar way to those without DA. The particles tend to accumulate along LCS, which act as barriers to transport. The possible discontinuities do not seem to affect or create artifacts in the FSLE field or LCS. When comparing the FSLE fields computed for consecutive days, the transition between them is smooth. There is no significant difference when comparing the variation of LCS for two consecutive days with or without DA between them. Furthermore, the experiments performed here show how we can reconstruct some LCS present in the ocean state when assimilating observations. In particular, the use of HFR data for assimilation helps to recreate the LCS present in the NR and to correct the dynamics and the transport in the region, as was demonstrated with the advection of particles.
@@
The four different areas in the IC from which we continuously deployed particles were selected in terms of their geographical location in the IC to evaluate the zonal and meridional transports in the channel and the connectivity between the different regions. The study could also be complemented by analysing the trajectories of particles launched at different sites as the ones shown in this work. Besides, a further quantitative analysis would be desirable, although, establishing a metric for this kind of analysis is difficult and not extended in the literature. FSLE fields should not be compared point-wise, as little differences in the position of LCS could affect the results, leading to an erroneous interpretation. Developing a valid metric to quantify LCS differences remains as a future work.
@@
OSSEs are an important tool to explore the capabilities of a future observing system design. Strictly, in scientific terms, it is always good to have as many observations as possible. However, as resources are limited, synergies between observing and modeling communities are needed to benefit mutually, and observing systems should rigorously be designed to meet user requirements and respond to societal needs (Davidson et al., 2019). For an optimal design of the observing system expansion further considerations could be taken into account. For instance, different locations for the antennas may be examined. For a final design, the decision should be jointly based on the scientific, and on the logistic and economic assessment, that are out of the scope of this work.
@@ 5 Conclusions
The objective of this study was to evaluate the impact of setting up a couple of antennas to complement the currently operating ones in the IC. We analyzed the impact of this new observing system on the transport properties through the LCS computation. The effect of data assimilation on the reconstruction of the LCS and its impact on the spreading of the advected particles has been assessed in a Lagrangian framework. A series of OSSEs assimilating HFR data along with traditional observing sources (SLA, SST, Argo) is presented here. The study is a continuation of the work from (Hernandez-Lasheras et al., 2021), where an OSE that is here used as reference to validate the OSSE framework was performed. The assessment of the OSSE is consistent with that of the reference OSE and the framework is considered suitable for the design and evaluation a future observing system.
The impact of a potential extension of the actual HFR system in the IC has been assessed. The two new antennas would provide a full coverage of the surface currents in the IC and could help to improve the forecasting of the circulation in the region. In circumstances where the flow regime represented by the model disagrees with the observed one, a diminution of up to 19% of the error can be expected when assimilating the future system observations, compared to the actual ones.
@@
JHL and BM conceptualized the OSSE. JHL, AO and IHC conceptualized the Lagrangian assessment. AS developed some of the Lagrangian tools used in this work. JHL conducted the experiments, performed the analysis and wrote the manuscript with the help of AO. All authors contributed in the discussion and review of the manuscript.
@@ Acknowledgements
This work was mostly done while Jaime Hernández-Laheras and Baptiste Mourre were at the SOCIB modelling facility. This research has been supported by the the EU Horizon 2020 JERICO-NEXT (grant agreement no. 654410) and EuroSea (grant agreement no. 862626). Alejandro Orfila acknowledges financial support from Projects LAMARCA (PID2021-123352OB-C31) funded by MICIN/AEI/10.13039/501100011033/ FEDER, UE; Tech2Coast (TED2021-130949B-I00) funded by MCIN/AEI/10.13039/ 501100011033 and by EU “NextGenerationEU/ PRTR” and LIFE AdaptCalaMillor – LIFE21 GIC/ES/101074227. The authors are indebted to the Balearic Islands Coastal Observing System (SOCIB) for the HF-Radar data. The present research was carried out in the framework of the AEI accreditation “Maria de Maeztu Centre of Excellence” given to IMEDEA (CSIC-UIB) (CEX2021-001198).

### v2

@@ 4 Discussion
The experiments presented here apply an approach which has not been used before for the design and extension of a HFR system. The NR is validated to give realistic simulations and the innovation distribution of the pseudo-observations present a similar distribution to the real data. The results obtained in the OSSE framework in terms of error reduction and correlation increment are of the same order as the ones obtained in the previous OSE work. For surface currents, there is a significant difference between the innovations obtained in OSE and OSSE, however we have seen that this difference is not systematic and depends on the analyzed period (Figure 6). Furthermore, when assimilating HFR, the error reduction and the increment in the correlation are also of the same order as the one achieved in the OSE.
@@
We have evaluated the impact on the surface currents in a wider area than that covered by the antennas, taking advantage of the full ocean state knowledge provided by the OSSE framework. It draws attention that the assimilation of the generic observing sources alone cannot correct the circulation in the area in the OSSE. In the previous OSE the assimilation of generic sources led to a good improvement compared to the CR. Even though the model’s response to SLA, SST and Argo observations is very similar to that obtained with real observations, the circulation seems to have a highly ageostrophic component. Therefore, the correction of SSH and density fields is not able to correct the surface currents in this region. This enhances the need of high-resolution surface current observations in coastal areas where satellite-derived geostrophic currents tend to fail (Vignudelli et al., 2019).
@@
The extension of the HFR system seems to be useful under certain conditions. The new antennas provide a moderate effect when NR and CR simulations reproduce similar circulation patterns. However, when both simulations present different circulation patterns, especially in the western side of the channel, the availability of surface current observations over the full channel is key to improve the model circulation. This can be interpreted as real situations in which the model is unable to reproduce the dynamics observed by the HFR system.
@@
The observation error for the HFR has been considered the same in all the domain. However, the error could be expected to be reduced in areas covered by three antennas where a larger number of radial observations can be used to generate a total reconstructed velocities. This approach using a spatially variable error depending on the number of available radial measurements at each point has been explored by some authors (Vandenbulcke et al., 2017). The improvement of the observation error, including correlated errors, remains an aspect to be explored in future studies. Besides, the generation of all pseudo-observations could be made more realistic by incorporating a spatially correlated noise.
@@
We used a Lagrangian approach to evaluate the impact of data assimilation on the surface transport. The Lagrangian techniques, such as FSLE, have been increasingly used in the last years. However, the effects of computing FSLEs using a model field sequentially corrected through DA remains still poorly studied. As particles are advected, the DA simulation’s discontinuities might impact the trajectories of the particles and the following computation of FSLE. Here we showed that simulations using DA behave in a similar way to those without DA. The particles tend to accumulate along LCS, which act as barriers to transport. The possible discontinuities do not seem to affect or create artifacts in the FSLE field or LCS. When comparing the FSLE fields computed for consecutive days, the transition between them is smooth. There is no significant difference when comparing the variation of LCS for two consecutive days with or without DA. Furthermore, the experiments performed here show how we can reconstruct some LCS present in the ocean state when assimilating observations. In particular, the use of HFR data for assimilation helps to recreate the LCS present in the NR and to correct the dynamics and the transport in the region, as was demonstrated with the advection of particles.
@@
The four different areas in the IC from which we continuously deployed particles were selected in terms of their geographical location in the IC to evaluate the zonal and meridional transports in the channel and the connectivity between the different regions. The study could also be complemented by analysing the trajectories of particles launched at different sites as the ones shown in this work. Besides, a further quantitative analysis would be desirable, even if, establishing a metric for this kind of analysis is difficult and not extended in the literature. FSLE fields should not be compared point-wise, as little differences in the position of LCS could affect the results, leading to an erroneous interpretation. Developing a valid metric to quantify LCS differences remains a future work.
@@
OSSEs are an important tool to explore the capabilities of a future observing system design. Strictly, in scientific terms, it is always good to have as many observations as possible. However, as resources are limited, synergies between observing and modeling communities are needed for a mutual benefit, and observing systems should rigorously be designed to meet user requirements and respond to societal needs (Davidson et al., 2019). Further considerations could be taken into account for an optimal design of the observing system expansion. For instance, different locations for the antennas may be examined. For a final design, the decision should be jointly based on the scientific, and on the logistic and economic assessment, that are out of the scope of this work.
@@ 5 Conclusions
The objective of this study was to evaluate the impact of setting up a couple of antennas to complement the currently operating ones in the IC. We analyzed the impact of this new observing system on the transport properties through the LCS computation. The effect of data assimilation on the reconstruction of the LCS and its impact on the spreading of the advected particles has been assessed in a Lagrangian framework. A series of OSSEs assimilating HFR data along with traditional observing sources (SLA, SST, Argo) is presented here. The study complements the work from Hernandez-Lasheras et al. (2021), which provides a reference OSE used to validate the OSSE framework performed here. The assessment of the OSSE is consistent with that of the reference OSE and the framework is considered suitable for the design and evaluation of the impact of future observing systems.
The impact of a potential extension of the actual HFR system in the IC has been assessed. The two new antennas would provide a full coverage of the surface currents in the IC and could help to improve the forecasting of the circulation in the region. In circumstances where the flow regime represented by the model would differ from the observed one, a diminution of up to 19% of the error can be expected when assimilating the future system observations, compared to the present situation.
@@
JHL and BM conceptualized the OSSE. JHL, AO and IHC conceptualized the Lagrangian assessment. AS developed some of the Lagrangian tools used in this work. JHL conducted the experiments, performed the analysis and wrote the manuscript with the help of AO and BM. All authors contributed to the discussion and review of the manuscript.
@@ Acknowledgements
This work was mostly performed within the SOCIB Modelling and Forecasting facility. This research has been supported by the the EU Horizon 2020 JERICO-NEXT (grant agreement no. 654410) and EuroSea (grant agreement no. 862626) projects. Alejandro Orfila acknowledges financial support from project LAMARCA (PID2021-123352OB-C31) funded by MICIN/AEI/10.13039/ 501100011033/ FEDER, UE. The authors are indebted to the Balearic Islands Coastal Observing System (SOCIB) for the HF-Radar data and modelling infrastructure. The authors are grateful to Emma Reyes for useful discussions about the Ibiza Channel HF radar setup. The authors also thank AEMET, the Spanish Meteorological Agency, for providing the HIRLAM atmospheric model fields and the Copernicus Service for satellite and modelling products (https://marine.copernicus.eu/). The present research was carried out in the framework of the AEI accreditation “Maria de Maeztu Centre of Excellence” given to IMEDEA (CSIC-UIB) (CEX2021-001198).